6NaNO3 + Al2O3 → 2Al(NO3)3 + 3Na2O



If 82.1g NaNO3 reacts with 100g Al2O3, what is the limiting reactant?

1
Expert's answer
2022-04-22T04:26:04-0400

6NaNO3 + Al2O3 → 2Al(NO3)3 + 3Na2O

n(NaNO3) = m(NaNO3)/Mr(NaNO3) = 82.1 g / 84.995 g/mol = 0.966 mol

n(Al2O3) = m(Al2O3)/Mr(Al2O3) = 100 g / 101.96 g/mol = 0.981 mol

6 mol NaNO3 react with 1 mol Al2O3

0.966 mol react with x mol Al2O3

x = 0.966 / 6 = 0.161 mol

NaNO3 is limiting reactant